OpenStudy (anonymous):

Choose the point on the terminal side of -210°. (-1, 1) (-1, square root of three) (square root of three, -1) (negative square root of three, 1)

OpenStudy (anonymous):

@amistre64 @Hero @IMStuck

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Simba!! :O

zepdrix (zepdrix):

If we spin 210 degrees `closewise`, which quadrant do we end up in? A little more than half a full spin `clockwise`, where are we at?

OpenStudy (anonymous):

Either the 4th, 1st,or 2nd I think

OpenStudy (anonymous):

around 45 degrees in the 1st quad?

zepdrix (zepdrix):

Spinning clockwise 210 is the same as spinning counter-closewise 150. (360-210). 150 degrees, I guess that puts us in ... quadrant 2.

OpenStudy (anonymous):

isn't that in the 4th quad?

zepdrix (zepdrix):

I'm on my tablet, grr I can't draw it for you :c hard to explain in words...

OpenStudy (anonymous):

this is my unit circle http://learn.flvs.net/educator/common/MrsKenyon/Unitcircle.pdf

zepdrix (zepdrix):

We're in quadrant 2 at 150 degrees. Along the unit circle we get the coordinate \(\Large\rm \left(-\frac{\sqrt3}{2},~\frac{1}{2}\right)\) But we can see that not all of our options are on the unit circle. They might be on the terminal side of our angle, just not on the unit circle.

zepdrix (zepdrix):

AHHH class is starting :c I gotta go. Option 4, is along the terminal side. poop. sorry

OpenStudy (anonymous):

Thanks!